OpenStudy (anonymous):

-6g+7g-7=9 Answer and work

OpenStudy (mathstudent55):

First, step, combine like terms on the left side. Can you do that and show what you get?

OpenStudy (anonymous):

\[-6g+7g-7=9\]\[g-7=9\]\[g-7+7=9+7\]\[g=16\] To check:\[-6(16)+7(6)-7=9\]\[-96+112-7=9\]\[9=9\] CHECK